The demand and supply equations are Qd = 100 – 10P and Qs = 30 + 5P. Compute for the

Price equilibrium and quantity equilibrium.


1
Expert's answer
2022-02-20T16:19:26-0500

"Q_s= Q_d"

"100-10P= 30+5P"

"15P= 70"

"P= 4.67"


"Q*= 100- 10(\frac{70}{15})"

"= \frac{(1500-700)}{15}= \frac {800}{15}"

= "53.33"